# A Diffused Background from Axion-like Particles in the Microwave Sky

### v1

@@ 4.2 Halo Modelling of photon-ALP Resonant conversion Power Spectrum
$\mathrm{C_{\ell,1h}^{ax}=\int_{z_{min}}^{z_{max}}dz\frac{dV_{c}}{dz}\int_{M_{min}}^{M_{max}}dM\frac{dn(M,z)}{dM}|\alpha_{\ell}(M,z)|^{2}.}$ (4.11)
Here $\mathrm{\frac{dV_{c}}{dz}}$ is the differential comoving volume and $\alpha_{\ell}$ is the angular harmonic transform of fluctuations (refer to Appendix C) due to the photon-ALP resonant conversion in a single cluster [15, 10]. $\alpha_{\ell}$’s depend on the electron density and magnetic field profiles of the clusters. The inference of $\alpha_{\ell}$’s and its dependency on frequency, coupling constant, and cluster profiles is explained in Appendix A when ALPs of all masses in the range $10^{-15}-10^{-11}$ eV are assumed to be generated in galaxy clusters. The integrals take into account the total number of clusters of various masses at different redshifts which lie in the cluster mass range ($\mathrm{M_{min}}$ to $\mathrm{M_{max}}$) and the redshift range being considered ($\mathrm{z_{min}}$ to $\mathrm{z_{max}}$). This contribution to the ALP power spectrum will be present even if there is no clustering effect due to gravitational attraction.
The ALP signal will also be correlated with the signal at locations outside its cluster. This is taken into account by the two-halo term (where we use the limber approximation applicable at $\mathrm{l}>20$) given as:
$\mathrm{C_{\ell,2h}^{ax}=\int_{z_{min}}^{z_{max}}dz\frac{dV_{c}}{dz}P_{m}\left(k=\frac{\ell+1/2}{r(z)},z\right)\times\left[\int_{M_{min}}^{M_{max}}dM\frac{dn(M,z)}{dM}b(M,z)\alpha_{\ell}(M,z)\right]^{2}}.$ (4.12)
Here $\mathrm{r(z)}$ is the comoving distance at redshift $\mathrm{z}$. The two-point halo correlation function has been expressed in terms of the matter power spectrum as
$\mathrm{P_{h}(k,M_{1},M_{2},z)=b(M_{1},z)b(M_{2},z)P_{m}(k,z),}$ (4.13)
@@
There will be several clusters in a certain mass interval within a given redshift interval. We have considered random electron density and magnetic field profiles for the galaxy clusters and the median of their $|\alpha_{\ell}\mathrm{(M,z)}|^{2}$’s as the contribution of these clusters to the ALP background spectrum. In principle, these single cluster power spectrums ($|\alpha_{\ell}\mathrm{(M,z)}|^{2}$’s) depend on the evolution of clusters based on their masses and redshifts. This in turn would allow one to relate the cluster electron density and magnetic field profiles with their masses at different redshifts and model the single cluster contribution to the power spectrum well. The single cluster power spectrums ($|\alpha_{\ell}\mathrm{(M,z)}|^{2}$’s) will also depend on the range of ALP masses that are being considered to be produced in the cluster and the variation of coupling constants with the ALP masses. We consider the case where ALPs of all masses in the range $10^{-15}-10^{-11}$ eV are produced in the galaxy clusters with the corresponding probabilities if the resonant condition is satisfied ($\mathrm{m_{a}=m_{\gamma}}$). The coupling constant has been assumed to be uniform for all ALP masses in this range ($\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$).
@@
Figure 2: The CMB is smooth at low angular scales, but if photon-ALP resonant conversion takes place, it will lead to additional fluctuations in the CMB. The RMS fluctuations from this conversion for a coupling constant of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$, would be of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$. The observed power spectrum at high multipoles in the presence of ALPs will not follow the usual dampening as observed for the case of the CMB-only map (blue line), but will increase with multipoles (orange line).
@@
The two-halo term increases at low multipoles ($\ell$ = 20 to 100) as the correlation is high, accompanied by a large number of clusters at low redshifts. With increasing multipoles, it then decreases as the number of clusters starts decreasing significantly with redshift and the halo correlation also decreases. The two-halo term may contribute more than one-halo term for the low-multipole range ($\mathrm{20<\ell<200}$) when $\mathrm{M_{min}}$ is low, as the low masses at low redshifts contribute significantly to the halo-halo correlation. Hence, the ALP spectrum shape will depend on the strength of the signal at low multipoles, which will be determined by the number of clusters contributing and their individual contributions, which themselves will depend on the cluster masses and profiles.
@@
There will be clusters at very high redshifts and with the polarized photon travelling through high and low redshift clusters, it may get depolarized due to turbulence or stochastic effects. Hence, it is mostly the low redshift clusters (z $<3.5$), (accompanied by the fact that they are in large numbers) that mainly contribute to the polarized ALP diffused signal. The polarization information will be lost as well if the beam size of the instrument is higher than the angular scale of the cluster on the sky.
@@
The CMB is very smooth at very high multipoles ($\ell>4000$), with low spatial fluctuations. If there is photon-ALP conversion in unresolved clusters, it will create a diffused background of the ALP signal which will cause additional fluctuations in the CMB at low angular scales or high multipoles (see Fig.2). For a coupling constant of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$, the additional RMS fluctuations would be of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$ at 145 GHz. The effect on the observed power spectrum is seen at high multipoles ($\ell>4000$) where the CMB-only map shows a dampening of power with increase in multipoles, while the power increases in the case of the CMB+ALPs map.
@@
Figure 3: This figure shows the one-halo Poissonian (solid lines) and two-halo clustering (dashed lines) components of the ALP background spectrum for a coupling constant of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$ at 145 GHz. The CMB polarization spectrum $\mathrm{D_{\ell}^{EE}}$ is shown in a dash-dotted line and is independent of frequency. Here $\mathrm{g_{12}}$ is dimensionless defined as $\mathrm{g_{12}=g_{a\gamma}\times 10^{12}\,GeV}$. The grey lines represent the noise power spectrum $\mathrm{N_{\ell}}$ corresponding to the beam size for frequency bands in the range 140-150 GHz for various detectors.
@@
The one-halo and two-halo power spectrums at 145 GHz are shown in Fig.3. These spectrums correspond to a minimum redshift $\mathrm{z_{min}=1}$. The grey lines represent the noise power spectrum $\mathrm{N_{\ell}}$ corresponding to the 140 - 150 GHz band for various detectors (SO, CMB-S4 and CMB-HD) as these are the bands with higher sensitivities compared to the ones at higher and lower frequencies.
@@
The increase in one-halo at low multipoles sets the maximum angular scale up to which ALPs are formed. At high multipoles, the shape of the ALP power spectrum will be independent of the strength and scale almost as $\ell^{0}$ ($\mathrm{D_{\ell}}$ varies as $\ell^{2}$) following the one-halo contribution. Here we have considered a randomly uniform orientation of the magnetic field at various locations in a cluster. This gives the $\ell^{2}$ dependence to $\mathrm{D_{\ell}}^{\alpha\alpha}$ at high multipoles.
@@
In principle, the variation of individual $|\alpha_{\ell}|$’s at high multipoles will depend on the magnetic field orientation within the cluster. The individual variations due to these $|\alpha_{\ell}|$’s will be suppressed as many clusters are integrated along the line of sight. The polarization information for a photon travelling from high redshift clusters may get lost due to depolarization from turbulence and stochastic effects in galaxy clusters. Thus, the low redshift clusters will impact the shape of the background spectrum the most as these are the clusters for which the polarization signal can be resolved. The ALP background spectrum will be correlated with the synchrotron background spectrum from galaxy clusters.
@@
The lensed CMB polarization power spectrum $\mathrm{D_{\ell}^{EE}}$ is independent of the frequency of observation. Also, its dependence on multipoles is well-known from the correlation between temperature and polarization spectrums. Other effects that induce polarization in the CMB include reionization and the polarized SZ effect. The reionization power spectrum decreases with an increase in frequency and also weakens at high multipoles. The polarized SZ power spectrum increases with multipoles but its strength is weak with fluctuations of the order of $10^{-8}$ K [13, 30]. The ALP power spectrum strength on the other hand increases with frequency in the radio and microwave regimes of the electromagnetic spectrum. The two-halo term may dominate at low multipoles ($20020<\ell<$ 200), while the one-halo term dominates at high multipoles. The high multipoles can be used to probe the ALP signal from the damped CMB power spectrum. The spectrum takes into account the ALP signals generated in various mass clusters, integrated over different redshifts. The individual features of the signal are thus suppressed and the power spectrum does not show spikes or high oscillations. This is in contrast to the CMB at low multipoles. For a low $\mathrm{z_{min}}$, this characteristic may not hold as resolved and well-luminous clusters may be contributing to the background. For the mass range $10^{13}$ to $\mathrm{7\times 10^{15}\,M_{\odot}}$, redshift range 1 to 7, and a coupling constant $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$, the one and two-halo contributions are shown in Fig. 3. Here $\mathrm{g_{12}}$ is dimensionless and is defined as $\mathrm{g_{12}\equiv g_{a\gamma}\times 10^{12}\,GeV}$. The ALP power spectrum crosses the CMB power spectrum at a lower multipole value for higher frequencies and vice versa. These spatial and spectral variations with respect to other polarized signals can be used to probe the ALP background spectrum.
@@ 4.4 Sources of variation of ALP diffused spectrum due to cosmological factors
Variation with minimum redshift: With a lower minimum redshift ($\mathrm{z_{min}}$), the background power spectrum increases (Fig.4), while it decreases with a higher minimum redshift. It is mainly the low redshift clusters that contribute to the ALP background signal as these are large in number and have their polarization signals intact. For the high redshift clusters, depolarization of the photons may occur as a result of multiple scatterings. With the upcoming improved detectors, we will be able to resolve clusters at lower redshifts (z < 1). The bright clusters at low redshifts could also significantly affect the shape and strength of the power spectrum, especially at low multipoles. Their contribution to the background spectrum need not be considered and constraints from them can be obtained using the analysis for resolved clusters (explained in a separate work [15]).
@@
Figure 4: Variation of the ALP power spectrum (solid: one-halo, dashed: two-halo) for various minimum redshifts $\mathrm{z_{min}}$. The power spectrum strength decreases with an increase in the minimum redshift which is considered as the clusters at low redshifts contribute to the background spectrum. The CMB E-mode spectrum is also plotted for reference
@@
Figure 5: This figure shows the variation of the ALP background spectrum with the minimum mass considered for the cluster range ($\mathrm{M_{min}}$). The spectrum increases with a lower minimum mass. This is due to the presence of more low mass clutsers as is depicted by the halo mass function. These low mass clusters contribute significantly to the spectrum, because of our random modelling of galaxy cluster profiles.
@@
Variation with cluster masses: The strength of the power spectrum also depends on the range of masses being considered. For different halo mass ranges, the spectrum decreases with decreasing mass range (Fig.5). The higher cluster masses contribute to low redshifts. The low cluster masses are less resolved and may contribute even at higher redshifts. The decrease in background spectrum will depend on the contribution of various mass clusters to the ALP background spectrum, which can be analyzed by studying the relation between cluster masses and their electron densities. Since the one-halo power spectrum is about two orders in magnitude higher than the two-halo contribution at the relevant scales corresponding to high multipoles ($\ell$ > 4000), we will not be plotting the two-halo contributions in any of the upcoming plots. But its contribution in the calculations will be considered.
@@
Figure 6: This figure shows the variation of ALP background spectrum with the ALP coupling constant $\mathrm{g_{a\gamma}}$. The power spectrum increases as the coupling strength increases, scaling as $\mathrm{g_{a\gamma}^{4}}$ The CMB E-mode spectrum is also plotted for reference. Here $\mathrm{g_{12}\equiv g_{a\gamma}\times 10^{12}\,GeV}$.
@@
Variation with photon-ALP coupling constant and ALPs masses: The amplitude of the ALP signal is proportional to the square of the coupling constant $\mathrm{g_{a\gamma}}$ (Eq.3.4), so the power spectrum scales as $\mathrm{g^{4}_{a\gamma}}$ (see Fig. 6). This is expected as the higher the coupling, the higher will be the distortion signal.
@@
We have considered ALPs of all masses in the range $\mathrm{10^{-15}-10^{-11}\,eV}$ are being produced. The ALP background spectrum will also depend on the ALP masses that may exist in nature. If ALPs of masses only in a certain subrange of the mass range considered, exist in nature, the power spectrum shall decrease. This decrease will be greater if only high mass ALPs exist ($\mathrm{10^{-13}-10^{-11}\,eV}$), while lower if low mass ALPs exist ($\mathrm{10^{-15}-10^{-13}\,eV}$). This is because the single cluster distortion spectrums $\alpha_{\ell}$’s are generally higher for low mass ALPs which are formed in the outer regions of clusters with low electron density and high conversion probabilities (see Eq.3.4).
@@
For a variation in coupling constant with ALP mass, signals would be generated in spherical shells in a spherically symmetric cluster. These shells would be visible as different-sized disks, with a larger disk corresponding to low mass ALPs. The signals from these different disks will then need to be summed up to obtain the net ALP distortion signal.

### v3

@@ 4.2 Halo Modelling of photon-ALP Resonant conversion Power Spectrum
$C_{\ell,1h}^{\mathrm{ax}}=\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\mathrm{d}z\frac{\mathrm{d}V_{c}}{\mathrm{d}z}\int_{M_{\mathrm{min}}}^{M_{\mathrm{max}}}\mathrm{d}M\frac{\mathrm{d}n(M,z)}{\mathrm{d}M}|\alpha_{\ell}(M,z)|^{2}.$ (4.11)
Here $\frac{\mathrm{d}V_{c}}{\mathrm{d}z}$ is the differential comoving volume and $\alpha_{\ell}$ is the angular harmonic transform of fluctuations (refer to Appendix C) due to the photon-ALP resonant conversion in a single cluster [34, 28]. $\alpha_{\ell}$’s depend on the electron density and magnetic field profiles of the clusters. The inference of $\alpha_{\ell}$’s and its dependency on frequency, coupling constant, and cluster profiles is explained in Appendix A when ALPs of all masses in the range $10^{-15}-10^{-11}$ eV are assumed to be generated in galaxy clusters. The integrals take into account the total number of clusters of various masses at different redshifts which lie in the cluster mass range ($M_{\mathrm{min}}$ to $M_{\mathrm{max}}$) and the redshift range being considered ($z_{\mathrm{min}}$ to $z_{\mathrm{max}}$). This contribution to the ALP power spectrum will be present even if there is no clustering effect due to gravitational attraction.
The ALP signal will also be correlated with the signal at locations outside its cluster. This is taken into account by the two-halo term (where we use the limber approximation applicable at $\ell>20$) given as:
$C_{\ell,2h}^{\mathrm{ax}}=\int_{z_{\mathrm{min}}}^{z_{\mathrm{max}}}\mathrm{d}z\frac{\mathrm{d}V_{c}}{\mathrm{d}z}P_{m}\left(k=\frac{\ell+1/2}{r(z)},z\right)\times\left[\int_{M_{\mathrm{min}}}^{M_{\mathrm{max}}}\mathrm{d}M\frac{\mathrm{d}n(M,z)}{\mathrm{d}M}b(M,z)\alpha_{\ell}(M,z)\right]^{2}.$ (4.12)
Here $r(z)$ is the comoving distance at redshift $z$. The two-point halo correlation function has been expressed in terms of the matter power spectrum as
$P_{h}(k,M_{1},M_{2},z)=b(M_{1},z)b(M_{2},z)P_{m}(k,z),$ (4.13)
@@
There will be several clusters in a certain mass interval within a given redshift interval. We have considered random electron density and magnetic field profiles for the galaxy clusters and the median of their $|\alpha_{\ell}(M,z)|^{2}$’s as the contribution of these clusters to the ALP background spectrum. In principle, these single cluster power spectra ($|\alpha_{\ell}(M,z)|^{2}$’s) depend on the evolution of clusters based on their masses and redshifts. This in turn would allow one to relate the cluster electron density and magnetic field profiles with their masses at different redshifts and model the single cluster contribution to the power spectrum well. The single cluster power spectra ($|\alpha_{\ell}(M,z)|^{2}$’s) will also depend on the range of ALP masses that are being considered to be produced in the cluster and the variation of coupling constants with the ALP masses. We consider the case where ALPs of all masses in the range $10^{-15}-10^{-11}$ eV are produced in the galaxy clusters with the corresponding probabilities if the resonant condition is satisfied ($m_{a}=m_{\gamma}$). The coupling constant has been assumed to be uniform for all ALP masses in this range ($g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$).
@@
Figure 2: The CMB is smooth at low angular scales, but if photon-ALP resonant conversion takes place, it will lead to additional fluctuations in the CMB. The RMS fluctuations from this conversion for a coupling constant of $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$, would be of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$. The observed power spectrum at high multipoles in the presence of ALPs will not follow the usual dampening as observed for the case of the CMB-only map (blue line), but will increase with multipoles (orange line).
@@
The two-halo term increases at low multipoles ($\ell$ = 20 to 100) as the correlation is high, accompanied by a large number of clusters at low redshifts. With increasing multipoles, it then decreases as the number of clusters starts decreasing significantly with redshift and the halo correlation also decreases. The two-halo term may contribute more than one-halo term for the low-multipole range ($\mathrm{20<\ell<200}$) when $M_{\mathrm{min}}$ is low, as the low masses at low redshifts contribute significantly to the halo-halo correlation. Hence, the ALP spectrum shape will depend on the strength of the signal at low multipoles, which will be determined by the number of clusters contributing and their individual contributions, which themselves will depend on the cluster masses and profiles.
@@
There will be clusters at very high redshifts and with the polarized photon travelling through high and low redshift clusters, it may get depolarized due to turbulence or stochastic effects. Hence, it is mostly the low redshift clusters ($z<3.5$), (accompanied by the fact that they are in large numbers) that mainly contribute to the polarized ALP diffused signal. The polarization information will be lost as well if the beam size of the instrument is higher than the angular scale of the cluster in the sky.
@@
The CMB is very smooth at very high multipoles ($\ell>4000$), with low spatial fluctuations. If there is photon-ALP conversion in unresolved clusters, it will create a diffused background of the ALP signal which will cause additional fluctuations in the CMB at low angular scales or high multipoles (see Fig.2). For a coupling constant of $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$, the additional RMS fluctuations would be of the order of $\mathrm{7.5\times 10^{-2}\,\mu K}$ at 145 GHz. The effect on the observed power spectrum is seen at high multipoles ($\ell>4000$) where the CMB-only map shows a dampening of power with increase in multipoles, while the power increases in the case of the CMB $+$ ALPs map.
@@
Figure 3: This figure shows the one-halo Poissonian (solid lines) and two-halo clustering (dashed lines) components of the ALP background spectrum for a coupling constant of $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$ at 145 GHz. The CMB polarization spectrum $D_{\ell}^{EE}$ is shown in a dash-dotted line and is independent of frequency. Here $g_{12}$ is dimensionless defined as $g_{12}=g_{a\gamma}\times 10^{12}\,\mathrm{GeV}$. The grey lines represent the noise power spectrum $N_{\ell}$ corresponding to the beam size for frequency bands in the range 140-150 GHz for various detectors.
@@
The one-halo and two-halo power spectra at 145 GHz are shown in Fig.3. These spectra correspond to a minimum redshift $z_{\mathrm{min}}=1$. The grey lines represent the noise power spectrum $N_{\ell}$ corresponding to the 140 - 150 GHz band for various detectors (SO, CMB-S4, and CMB-HD) as these are the bands with higher sensitivities compared to the ones at higher and lower frequencies.
@@
The increase in one-halo at low multipoles sets the maximum angular scale up to which ALPs are formed. At high multipoles, the shape of the ALP power spectrum will be independent of the strength and scale almost as $\ell^{0}$ ($D_{\ell}$ varies as $\ell^{2}$) following the one-halo contribution. Here we have considered a randomly uniform orientation of the magnetic field at various locations in a cluster. This gives the $\ell^{2}$ dependence to $D_{\ell}^{\alpha\alpha}$ at high multipoles.
@@
In principle, the variation of individual $|\alpha_{\ell}|$’s at high multipoles will depend on the magnetic field orientation within the cluster. The individual variations due to these $|\alpha_{\ell}|$’s will be suppressed as many clusters are integrated along the line of sight. The polarization information for a photon travelling from high redshift clusters may get lost due to depolarization from turbulence and stochastic effects in galaxy clusters. Thus, the low redshift clusters will impact the shape of the background spectrum the most as these are the clusters for which the polarization signal can be resolved. The ALP background spectrum will be correlated with the synchrotron background spectrum from galaxy clusters. In this analysis, we have also considered that the coherent length scale of magnetic fields is larger than the beam size of the CMB experiments for it to be detectable. For coherent scales of magnetic field that are smaller than the beam size, the polarised signal will be smeared out. We have taken this effect into account in our analysis. Thus a high-resolution beam is important, so that the polarization information is well preserved. The polarized distortion information from high redshift clusters will be lost as the angular scale of the magnetic field will be smaller than the beam size. Thus for the upcoming CMB experiments, it is mostly the low redshift clusters that will significantly contribute to the background spectrum. With high angular resolution CMB experiments in the future, measurement of the signal from the high redshift clusters will improve by using multiple frequency channels. Efficient foreground cleaning techniques [49, 50, 51] on the polarization maps will be useful in reconstructing this signal from data in the future. In addition to the smearing due to the instrument beam, there can be smearing along the line of sight if there are multiple galaxy clusters along the same line of sight where resonant photon-ALPs conversion can take place. However, this effect is low as the probability of multiple galaxy clusters along the same line of sight is less. Also, the amplitude of the diffused spectrum will increase with the strength of the magnetic fields for these low redshift clusters.
@@
The lensed CMB polarization power spectrum $D_{\ell}^{EE}$ is independent of the frequency of observation. Also, its dependence on multipoles is well-known from the correlation between temperature and polarization spectra. Other effects that induce polarization in the CMB include reionization and the polarized SZ effect. The reionization power spectrum decreases with an increase in frequency and also weakens at high multipoles. The polarized SZ power spectrum increases with multipoles but its strength is weak with fluctuations of the order of $10^{-8}$ K [32, 52]. The ALP power spectrum strength on the other hand increases with frequency in the radio and microwave regimes of the electromagnetic spectrum. The two-halo term may dominate at low multipoles ($20020<\ell<$ 200), while the one-halo term dominates at high multipoles. The high multipoles can be used to probe the ALP signal from the damped CMB power spectrum. The spectrum takes into account the ALP signals generated in various mass clusters, integrated over different redshifts. The individual features of the signal are thus suppressed and the power spectrum does not show spikes or high oscillations. This is in contrast to the CMB at low multipoles. For a low $z_{\mathrm{min}}$, this characteristic may not hold as resolved, and well-luminous clusters may be contributing to the background. For the mass range $10^{13}$ to $\mathrm{7\times 10^{15}\,M_{\odot}}$, redshift range 1 to 7, and a coupling constant $g_{a\gamma}=10^{-12}\,\mathrm{GeV^{-1}}$, the one and two-halo contributions are shown in Fig. 3. Here $g_{12}$ is dimensionless and is defined as $g_{12}\equiv g_{a\gamma}\times 10^{12}\,\mathrm{GeV}$. The ALP power spectrum crosses the CMB power spectrum at a lower multipole value for higher frequencies and vice versa. These spatial and spectral variations with respect to other polarized signals can be used to probe the ALP background spectrum. Also, the diffused ALP signal can potentially impact the study of small-scale B-mode polarization signals, which are mainly attributed to gravitational lensing of CMB photons due to large-scale structure [53, 54]. However, as the spectral shape and the angular power spectrum due to the ALP distortion are different from CMB B-mode polarization, it is possible to distinguish the two using multiple frequency bands. Furthermore, a joint estimation of the two signals will be possible as well from the future detectors.
@@ 4.4 Sources of variation of ALP diffused spectrum due to cosmological factors
Variation with minimum redshift: With a lower minimum redshift ($z_{\mathrm{min}}$), the background power spectrum increases (Fig.4), while it decreases with a higher minimum redshift. It is mainly the low redshift clusters that contribute to the ALP background signal as these are large in number and have their polarization signals intact. For the high redshift clusters, depolarization of the photons may occur as a result of multiple scatterings. With the upcoming improved detectors, we will be able to resolve clusters at lower redshifts ($z<1$). The bright clusters at low redshifts could also significantly affect the shape and strength of the power spectrum, especially at low multipoles. Their contribution to the background spectrum need not be considered and constraints from them can be obtained using the analysis for resolved clusters (explained in a separate work [34]).
@@
Figure 4: Variation of the ALP power spectrum (solid: one-halo, dashed: two-halo) for various minimum redshifts $z_{\mathrm{min}}$. The power spectrum strength decreases with an increase in the minimum redshift which is considered as the clusters at low redshifts contribute to the background spectrum. The CMB E-mode spectrum is also plotted for reference
@@
Figure 5: This figure shows the variation of the ALP background spectrum with the minimum mass considered for the cluster range ($M_{\mathrm{min}}$). The spectrum increases with a lower minimum mass. This is due to the presence of more low mass clusters as is depicted by the halo mass function. These low mass clusters contribute significantly to the spectrum, because of our random modelling of galaxy cluster profiles.
@@
Variation with cluster masses: The strength of the power spectrum also depends on the range of masses being considered. For different halo mass ranges, the spectrum decreases with decreasing mass range (Fig.5). The higher cluster masses contribute to low redshifts. The low cluster masses are less resolved and may contribute even at higher redshifts. The decrease in background spectrum will depend on the contribution of various mass clusters to the ALP background spectrum, which can be analyzed by studying the relation between cluster masses and their electron densities. Since the one-halo power spectrum is about two orders in magnitude higher than the two-halo contribution at the relevant scales corresponding to high multipoles ($\ell$ > 4000), we will not be plotting the two-halo contributions in any of the upcoming plots. But its contribution to the calculations will be considered.
@@
Figure 6: This figure shows the variation of ALP background spectrum with the ALP coupling constant $g_{a\gamma}$. The power spectrum increases as the coupling strength increases, scaling as $g_{a\gamma}^{4}$ The CMB E-mode spectrum is also plotted for reference. Here $g_{12}\equiv g_{a\gamma}\times 10^{12}\,\mathrm{GeV}$.
@@
Variation with photon-ALP coupling constant and ALPs masses: The amplitude of the ALP signal is proportional to the square of the coupling constant $g_{a\gamma}$ (Eq.3.5), so the power spectrum scales as $g^{4}_{a\gamma}$ (see Fig. 6). This is expected as the higher the coupling, the higher will be the distortion signal.
@@
We have considered ALPs of all masses in the range $\mathrm{10^{-15}-10^{-11}\,eV}$ are being produced. The ALP background spectrum will also depend on the ALP masses that may exist in nature. If ALPs of masses only in a certain subrange of the mass range considered, exist in nature, the power spectrum shall decrease. This decrease will be greater if only high mass ALPs exist ($\mathrm{10^{-13}-10^{-11}\,eV}$), while lower if low mass ALPs exist ($\mathrm{10^{-15}-10^{-13}\,eV}$). This is because the single cluster distortion spectra $\alpha_{\ell}$’s are generally higher for low mass ALPs which are formed in the outer regions of clusters with low electron density and high conversion probabilities (see Eq.3.5).
If ALPs of only a particular mass exist, the resonant conversion can occur only at two locations along a line of sight. However, if ALPs of different masses co-exist, the depolarization of the polarization components of the signal can occur when the magnetic field orientation varies, and the different mass ALPs are formed along the same line of sight. This is the case for the overlapping region of the different mass ALP signal disks and lines of sight near the cluster center. But the intensity of the signal will increase in the overlapping region.
@@
For a variation in coupling constant with ALP mass, signals would be generated in spherical shells in a spherically symmetric cluster. These shells would be visible as different-sized disks, with a larger disk corresponding to low mass ALPs. The intensities from these different disks will then need to be summed up to obtain the net ALP distortion signal.